OpenStudy (anonymous):

Simplify if possible. If an expression cannot be simplified write not possible. 2 square root 3 + square root 27- square root 1/3 plz help!!…

jhonyy9 (jhonyy9):

2sqrt3 +sqrt27 -sqrt(1/3) =2sqrt3 +sqrt(9*3) -(sqrt3)/3 = =2sqrt3 +3sqrt3 -(sqrt3)/3 = 5sqrt3 -(sqrt3)/3 =(15sqrt3 -sqrt3)/3 =(14sqrt3)/3